OpenStudy (anonymous):

x2-9x+20 is it (x-5) (x-4)?

OpenStudy (anonymous):

yes, it comes after factorisation

OpenStudy (anonymous):

YA EXACTLY

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Yeah, x * x = x^2 x * -4 = -4x -5 * x = -5x -5 * -4 = 20 x^2 - 4x - 5x + 20 x^2 - 9x + 20